D. Read and analyze the problem below. Solve and show complete solution.


Two stations, located at M(1.5, 0) and N(1.5, 0) (units are in km), simultaneously send sound

signals to a ship, with the signal traveling at the speed of 0.33 km/s. If the signal from N was

received by the ship four seconds before the signal it received from M, find the equation of the

curve containing the possible location of the ship.



1
Expert's answer
2021-03-18T04:09:23-0400

"\text {let us choose a coordinate system so that the X axis}"

"\text{ passes through the stations }M,N"

"\text{and the Y axis through the middle of the segment} MN"

"\text {let } MN= 2c\text{ then}"

"M(-c,0);N(-c;0)"

"S(x,y) - \text{coordinates of points satisfying the condition:}"

"SM - SN = 2a"

"2a -\text {this is the distance traveled by the sound for the difference in time}"

"\text{of signal fixation by the receiving stations}M,N"

"SM =\sqrt{(x+c)^2+y^2}"

"SN =\sqrt{(x-c)^2+y^2}"

"SM>SN"

"\sqrt{(x+c)^2+y^2}>\sqrt{(x-c)^2+y^2}"

"(x+c)^2+y^2>(x-c)^2+y^2"

"(x+c)^2>(x-c)^2"

"x^2+2xc+c^2>x^2-2xc+c^2"

"4xc>0"

"x>0(1)"

"SM - SN = 2a"

"\sqrt{(x+c)^2+y^2} - \sqrt{(x-c)^2+y^2} =2a(2)"

"\sqrt{(x+c)^2+y^2} =2a+\sqrt{(x-c)^2+y^2}"

"(x+c)^2+y^2 =4a^2+4a\sqrt{(x-c)^2+y^2}+(x-c)^2+y^2"

"-4a^2+4xc =4a\sqrt{(x-c)^2+y^2}"

"-a^2+xc =a\sqrt{(x-c)^2+y^2}"

"a^4-4axc+x^2c^2=a^2((x-c)^2+y^2)"

"x^2(c^2-a^2)-a^2y^2=a^2(c^2-a^2)"

"\frac{x^2}{a^2}-\frac{y^2}{c^2-a^2}=1(3)"


"\text{Combining formula (1) and (2), we got the total solution}:"

"S(x,y)"

"\frac{x^2}{a^2}-\frac{y^2}{c^2-a^2}=1\text{ for}x>0"


"\text{This solution can be obtained in a shorter way using the definition of hyperbola}"

"\text{let's apply specific data from the task:}"

"M(1.5,0);N(1.5,0)"

"c=0"

"M(0,0);N(0,0)"

"2a= 0.33*4=1.32"

"\text{According to the formula (2)}"

"\sqrt{(x+c)^2+y^2} - \sqrt{(x-c)^2+y^2} =2a"

"\sqrt{(x+0)^2+y^2} - \sqrt{(x-0)^2+y^2} =1.32"

"0=1.32\text{ its false}"

"\text{For the given initial conditions, the problem has no solution}"

Answer:For the given initial conditions, the problem has no solution